OpenStudy (anonymous):

3b-9-8b=11 solve the equation. show work check answer

OpenStudy (anonymous):

I just showed you how to do the same thing before! Have a try! All you have to do is group the like terms together until you have b on the left and a numerical value on the right.

OpenStudy (anonymous):

so wait subtract 8b from 3b? then drop down the answer and the 9 and what it eaquals to then subtract nine from both sides

OpenStudy (anonymous):

correct so step it out and I'll check it for you

OpenStudy (anonymous):

3b-9-8b=11 5b-9=11 +9 +9 ----------- 14b=18 14/14=0 18/14=

OpenStudy (anonymous):

Not correct but nice try! 3b-8b=-5b not +5b Also when you add 9 to both sides, 9 cant be added to 5b as it has a b in it. so +9 on the left hand side just deletes the -9 (as -9+9=0) +9 on the right hand side is 11+9=20 so go back through and see how you go

OpenStudy (anonymous):

wait where do i start?

OpenStudy (anonymous):

3b-9-8b=11 collect like terms first on the left hand side, then just try and get b on left side and numeric value on right side following the rules I gave you above. You were on the right track before, you just werent calculating correctly.

OpenStudy (anonymous):

3b-9-8b=11 8b-3b= -5b -5b-9=11 -5b/5=0 11/5=2.2

OpenStudy (anonymous):

No your getting confused look how I step it out and you should get what I mean 3b-9-8b=11 ok first lets group the b values together 3b-8b-9=11 -5b-9=11 Then I need to get rid of the -9 on the left hand side so I +9 to both sides. Remember you can only +9 to numeric values, not any with a b term in it. -5b-9(+9)=11(+9) -5b=20 Then I need to get the -5 over to the other side, so I divide both sides by -5 -5b/5=-20/5 b=-4

OpenStudy (anonymous):

how do i check this one

OpenStudy (anonymous):

????

OpenStudy (anonymous):

same as before just replace -4 for x in the original problem

OpenStudy (anonymous):

sorry for b

OpenStudy (anonymous):

???

OpenStudy (anonymous):

what

OpenStudy (anonymous):

did you try checking the equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

work out for you?